Harry Potter ~ Many people love reading books like Harry Potter. Historically, it has been estimated that the proportion of book readers, who like to read fiction, is 0.478. Camille is a Harry Potter fan and a fan of fiction books. She wants to determine if the actual proportion of book readers who like to read fiction is different from the historical value. From a recent survey of randomly selected 14,999 book readers, she learns that 7,049 like to read fiction.

A normal distribution is used to carry out the following hypothesis test:

H0:p=0.478 vs  HA:p≠0.478.

What is the null standard error used to calculate the test statistic for this hypothesis test? Give your answer to 4 decimal places.
Answer #1

Given that,

p = 0.478

1 - p = 1-0.478=0.522

n = 14999

standard error SE =  [P0 ( 1 - P0 ) / n] =   [(0.478*0.522) / 14999 ] = 0.0041

standard error SE = 0.0041